OpenStudy (anonymous):

Solving Systems of Equations: 4x=3y-2 and 18=3x+y...How do I solve this???

OpenStudy (anonymous):

18 = 3x + y y = 18 - 3x Now plug it in instead of y in the other equation and solve for x.

OpenStudy (anonymous):

Ok,I got x=52/11.

OpenStudy (anonymous):

Check you answer, x = 4.

OpenStudy (anonymous):

Can you please explain how to get that answer?

OpenStudy (anonymous):

4x = 3y - 2 I plug in y = 18 - 3x 4x = 3(18 - 3x) -2 4x = 54 - 9x -2 13x = 52 x = 4

OpenStudy (anonymous):

Now replace x by 4 and find y.

OpenStudy (anonymous):

Thank you :)